OpenStudy (lovelycharm):

If sin Θ = 5 over 6, what are the values of cos Θ and tan Θ?

hartnn (hartnn):

what have you tried?

OpenStudy (lovelycharm):

wym

hartnn (hartnn):

there are several methods to solve this, triangle method, trigonometric identities, etc... which one are you comfortable with?

OpenStudy (lovelycharm):

triangle

hartnn (hartnn):

\(\sin^2 x + \cos^2 x =1\) you have sin value, plug that in and find the cos value!

OpenStudy (lovelycharm):

where did u get those number from

hartnn (hartnn):

thats a trigonometric identity that you can use. sin x = 5/6 is given so \((5/6)^2 + \cos^2 x =1\) only cos x is unknown here, isolate it and find its value :)

OpenStudy (lovelycharm):

cos will be 11

hartnn (hartnn):

how did u get that? can you show me your workings?

OpenStudy (lovelycharm):

i added 6+5

hartnn (hartnn):

let me show you the workings and you let me know if you have any doubts in that, \((5/6)^2 +\cos^2 x =1 \\ \cos^2 x = 1 -(5/6)^2 \\ \cos^2 x = 1- 25/36 \\ \cos^2 x = (36-25)/36 \\ \cos^2 x = 11/36 \\ \cos x = \sqrt{11/36} \\ \cos x = \dfrac{\sqrt {11}}{6} \\ \) thats your answer for cos. tan is very easy, just tan x = sin x/ cos x try to get it :)

OpenStudy (lovelycharm):

thank u